OpenStudy (anonymous):

How many ounces of a 22% alcohol solution must be mixed with 23 ounces of a 5% alcohol solution to produce a 10% alcohol solution? Round to the nearest tenth of an ounce. Can anybody explain how to do this or the other question I've posted?

OpenStudy (anonymous):

yes i think i can. you add x ounces it will have .22x ounces of alcohol . the 23 once solution contains .05*23 = 1.15 onces of alcohol. when you add x ounces your total will be 23 + x and you want that to be ten percent alcohol, which will be .10(23+x)

OpenStudy (anonymous):

set them equal to solve:\[.22x+1.15=.10(23+x)\] \[.22x+1.15=2.3+.10x\] \[.22x-.10x=2.3-1.15\] \[.12x=1.15\] \[x=\frac{1.15}{.12}=9.58\] rounded